OpenStudy (anonymous):

Can anyone tell me the formula for the surface are of a cylinder. I can do the math, I just need the steps.

OpenStudy (anonymous):

S = 2*pi*r(h+r)

OpenStudy (anonymous):

s = area, r = radius, h = height

OpenStudy (anonymous):

you use the radius twice?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

okay. Thanks!

OpenStudy (anonymous):

How do you get this formula: The base area = pi*r² The lateral area = 2*pi*r*h The surface area = 2*base area + lateral area, right? So, 2(pir²) + 2pirh = 2pir(h+r) :))